# C-Pack: Packed Resources For General Chinese Embeddings

### v3

@@ Abstract.
We introduce C-Pack, a package of resources that significantly advance the field of general text embeddings for Chinese. C-Pack includes three critical resources. 1) C-MTP is a massive training dataset for text embedding, which is based on the curation of vast unlabeled corpora and the integration of high-quality labeled corpora. 2) C-MTEB is a comprehensive benchmark for Chinese text embeddings covering 6 tasks and 35 datasets. 3) BGE is a family of embedding models covering multiple sizes. Our models outperform all prior Chinese text embeddings on C-MTEB by more than +10% upon the time of the release. We also integrate and optimize the entire suite of training methods for BGE. Along with our resources on general Chinese embedding, we release our data and models for English text embeddings. The English models also achieve state-of-the-art performance on the MTEB benchmark; meanwhile, our released English data is 2 times larger than the Chinese data. Both Chinese and English datasets are the largest public release of training data for text embeddings. All these resources are made publicly available at https://github.com/FlagOpen/FlagEmbedding.
@@
journalyear: 2024copyright: acmlicensedconference: Proceedings of the 47th International ACM SIGIR Conference on Research and Development in Information Retrieval; July 14–18, 2024; Washington, DC, USA.booktitle: Proceedings of the 47th International ACM SIGIR Conference on Research and Development in Information Retrieval (SIGIR ’24), July 14–18, 2024, Washington, DC, USAisbn: 979-8-4007-0431-4/24/07doi: 10.1145/XXXXXX.XXXXXXccs: Information systems Retrieval models and ranking
@@ 1. Introduction
Text embedding is a long-standing topic in natural language processing and information retrieval. By representing texts with latent semantic vectors, text embedding can support various applications, e.g., web search, question answering, and retrieval-augmented language modeling (Karpukhin et al., 2020; Lewis et al., 2020; Guu et al., 2020; Xiong et al., 2020). The recent popularity of large language models (LLMs) has made text embeddings even more important. Due to the inherent limitations of LLMs, such as world knowledge and action space, external support via knowledge bases or tool use is necessary. Text embeddings are critical to connect LLMs with these external modules (Borgeaud et al., 2022; Qin et al., 2023).
@@
The remaining part of this paper is organized as follows. We discuss the related works about general text embeddings in Section 2. We present a detailed introduction about C-Pack resources in Section 3. We make comprehensive empirical analysis for the value of C-Pack in Section 4. Finally, we conclude our work in Section 5.
@@ 2. Related Work
The importance of general text embedding is widely recognized, not only for its wide usage in typical applications, like web search and question answering (Karpukhin et al., 2020) but also due to its fundamental role in augmenting large language models (Lewis et al., 2020; Guu et al., 2020; Borgeaud et al., 2022; Izacard et al., 2022; Shi et al., 2023). Compared with the conventional task-specific methods, the general text embedding needs to be extensively applicable in different scenarios. In recent years, there has been a continual effort in this field, where a series of well-known works are proposed, like Contriever (Izacard et al., 2021), GTR (Ni et al., 2021b), sentence-T5 (Ni et al., 2021a), Sentence-Transformer (Reimers and Gurevych, 2019), E5 (Wang et al., 2022a), OpenAI text embedding (Neelakantan et al., 2022), etc. Although it remains an open problem, recent studies highlight the following important factors.
@@ 4.2. Detailed Analysis
5. Conclusion
In this work, we present C-Pack to advance the progress of general Chinese embedding. C-Pack is made up of a comprehensive package of critical resources: 1) The benchmark C-MTEB, covering 6 major tasks of embeddings and 35 datasets, making it the most comprehensive benchmark to evaluate the generality of Chinese text embeddings. 2) The training data C-MTP, created by the elaborate and comprehensive curation of massive unlabeled corpora and the integration of high-quality labeled data. Its unprecedented scale, diversity, and quality contribute to the superior generality of the text embedding models. 3) The models BGE, which are empirically competitive for both direct usage and downstream fine-tuning. The different model sizes provide people with the flexibility to trade off efficiency and embedding quality. 4) The entire training recipe is also provided along with these resources. The public release of C-Pack enjoys a widespread recognition from the technique communities. It substantially facilitates the application of general Chinese embedding and also paves the way for its future development.

### v4

@@ Abstract.
We introduce C-Pack, a package of resources that significantly advances the field of general text embeddings for Chinese. C-Pack includes three critical resources. 1) C-MTP is a massive training dataset for text embedding, which is based on the curation of vast unlabeled corpora and the integration of high-quality labeled corpora. 2) C-MTEB is a comprehensive benchmark for Chinese text embeddings covering 6 tasks and 35 datasets. 3) BGE is a family of embedding models covering multiple sizes. Our models outperform all prior Chinese text embeddings on C-MTEB by more than +10% upon the time of the release. We also integrate and optimize the entire suite of training methods for BGE. Along with our resources on general Chinese embedding, we release our data and models for English text embeddings. The English models also achieve state-of-the-art performance on the MTEB benchmark; meanwhile, our released English data is 2 times larger than the Chinese data. Both Chinese and English datasets are the largest public release of training data for text embeddings. All these resources are made publicly available at https://github.com/FlagOpen/FlagEmbedding.
@@
journalyear: 2024copyright: acmlicensedconference: Proceedings of the 47th International ACM SIGIR Conference on Research and Development in Information Retrieval; July 14–18, 2024; Washington, DC, USAbooktitle: Proceedings of the 47th International ACM SIGIR Conference on Research and Development in Information Retrieval (SIGIR ’24), July 14–18, 2024, Washington, DC, USAdoi: 10.1145/3626772.3657878isbn: 979-8-4007-0431-4/24/07ccs: Information systems Retrieval models and ranking
@@ 1. Introduction
Text embedding is a long-standing topic in natural language processing and information retrieval. By representing texts with latent semantic vectors, text embedding can support various applications, e.g., web search, question answering, and retrieval-augmented language modeling (Zhang et al., 2022; Xiao et al., 2021; Lewis et al., 2020; Xiao et al., 2022a). The recent popularity of large language models (LLMs) has made text embeddings even more important. Due to the inherent limitations of LLMs, such as world knowledge and action space, external support via knowledge bases or tool use is necessary. Text embeddings are critical to connect LLMs with these external modules (Borgeaud et al., 2022; Qin et al., 2023).
@@
The remaining part of this paper is organized as follows. We discuss the related works about general text embeddings in Section 2. We present a detailed introduction about C-Pack resources in Section 3. We make comprehensive empirical analysis for the value of C-Pack in Section 4. Finally, we conclude our work in Section LABEL:sec:conclude.
@@ 2. Related Work
The importance of general text embedding is widely recognized, not only for its wide usage in typical applications, like web search and question answering (Xiao et al., 2022b; Karpukhin et al., 2020) but also due to its fundamental role in augmenting large language models (Lewis et al., 2020; Guu et al., 2020; Borgeaud et al., 2022; Izacard et al., 2022; Shi et al., 2023). Compared with the conventional task-specific methods, the general text embedding needs to be extensively applicable in different scenarios. In recent years, there has been a continual effort in this field, where a series of well-known works are proposed, like Contriever (Izacard et al., 2021), GTR (Ni et al., 2021b), sentence-T5 (Ni et al., 2021a), Sentence-Transformer (Reimers and Gurevych, 2019), E5 (Wang et al., 2022a), OpenAI text embedding (Neelakantan et al., 2022), etc. Although it remains an open problem, recent studies highlight the following important factors.